OpenStudy (anonymous):

Write as an algebraic expression in x free of trigonometric or inverse trigonometric functions: cos(cot^-1x)

OpenStudy (anonymous):

1/Sqrt[1 + 1/x^2]

OpenStudy (noelgreco):

 So the answer should be: \[\cos (\cot ^{-1}x)= \frac{ x }{ \sqrt{x ^{2}+1} }\]